# MAGPIE: Multilevel–Adaptive–Guided solver for ptychographic phase retrieval

### v4

@@ Appendix D Useful Lemmas
$\widehat{\bm{z}_{H,k}}=\bm{z}^{\prime}_{H,k}+\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}_{H}\bm{z}^{\prime}_{H,k}\right)\,,$
@@
$\bm{e}_{H,k}=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\bm{Q}_{H}\bm{z}_{H,k}^{\prime}-\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\right)\,.$
Then, by Lemma D.1 and Lemma D.2, we have
$\bm{e}_{H,k}$ $=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\bm{Q}_{H}\bm{z}_{H,k}^{\prime}-\bm{I}_{h}^{H}\left(\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\bm{W}_{\bm{z}}}{\bm{Q}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,,$
$=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\left(\bm{Q}_{H}\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\bm{z}^{\prime}_{k})-\bm{Q}_{H}\bm{I}_{h}^{H}\left(\frac{\bm{W}_{\bm{z}}}{\bm{Q}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\left(\bm{z}^{\prime}_{k}-\frac{1}{\bm{Q}}\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\bm{I}_{h}^{H}\left(\frac{\bm{W}_{\bm{z}}}{|\bm{Q}|^{2}}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\bm{I}_{h}^{H}\left(\frac{1}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,,$
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$
The proof follows from Proposition G.1 with
$\bm{A}=\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$
Appendix H Proof for Proposition 3.4
@@
$\bm{z}_{k}^{+}=\operatornamewithlimits{argmin}_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{\prime}\right|^{2}\,,$
@@
$\bm{e}_{k}=\bm{z}_{k}^{+}-\bm{z}_{k}^{\prime}=-\frac{1}{\bm{u}+|\bm{Q}|^{2}}\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\,,$
@@
$\bm{e}_{H,k}=-\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$
@@
$\bm{e}_{k}(\bm{g})=-\frac{1}{\bm{u}+|\bm{Q}|^{2}}\bm{g}\quad\text{and}\quad\bm{e}_{H,k}(\bm{g})=-\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\bm{g}\,.$
Since $\|\bm{I}_{h}^{H}\|_{\infty}=1$ by Lemma D.4, it is straightforward to see that
@@
$\max_{\|\bm{g}\|_{\infty}\leq 1}\left|\bm{e}_{H,k}(\bm{g})\right|_{r}$ $=\max_{\|\bm{g}_{H}\|_{\infty}\leq 1}\left|\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{g}_{H}\right|_{r}\,,$
$=\left(\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)_{r}\,,$
@@
$\max_{\|\bm{g}\|_{\infty}\leq 1}\left|\bm{I}_{h}^{H}\bm{e}_{k}(\bm{g})\right|_{r}=\left(\bm{I}_{h}^{H}\frac{1}{\bm{u}+|\bm{Q}|^{2}}\right)_{r}\,.$
Finally, using Lemma D.1, Lemma D.2, and Lemma D.3, we have:
$\bm{u}_{H}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\bm{u}\,,$
$\bm{u}_{H}+|\bm{Q}_{H}|^{2}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\bm{u}+|\bm{Q}_{H}|^{2}\,,$
$\bm{u}_{H}+|\bm{Q}_{H}|^{2}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{I}_{h}^{H}\left(\bm{u}+|\bm{Q}|^{2}\right)\,,$
$\bm{I}_{h}^{H}\left(\frac{1}{\bm{u}+|\bm{Q}|^{2}}\right)$ $=\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,,$
$\max_{\|\bm{g}_{2}\|_{\infty}\leq 1}\left|\bm{I}_{h}^{H}\bm{e}_{k}(\bm{g}_{2})\right|_{r}$ $=\max_{\|\bm{g}_{1}\|_{\infty}\leq 1}\left|\bm{e}_{H,k}(\bm{g}_{1})\right|_{r}\quad\text{for all }r\,.$

### v6

@@ Appendix E Useful Lemmas
$\widehat{\bm{z}_{H,k}}=\bm{z}^{\prime}_{H,k}+\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\left(\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}_{H}\odot\bm{z}^{\prime}_{H,k}\right)\,$,
@@
$\bm{e}_{H,k}=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\left(\bm{Q}_{H}\odot\bm{z}_{H,k}^{\prime}-\mathcal{R}_{H,k}\left(\bm{z}_{k}^{(j)}\right)\right)\,.$
Then, by Lemma E.1 and Lemma E.2, we have
$\bm{e}_{H,k}$ $=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\left(\bm{Q}_{H}\odot\bm{z}_{H,k}^{\prime}-\bm{I}_{h}^{H}\left(\frac{(\bm{I}_{H}^{h}\bm{Q}_{H})\odot\bm{W}_{\bm{z}}}{\bm{Q}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,$,
$=-\frac{\overline{\bm{Q}_{H}}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\left(\bm{Q}_{H}\odot\bm{I}_{h}^{H}(\bm{W}_{\bm{z}}\odot\bm{z}^{\prime}_{k})-\bm{Q}_{H}\odot\bm{I}_{h}^{H}\left(\frac{\bm{W}_{\bm{z}}}{\bm{Q}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,$,
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\bm{I}_{h}^{H}\left(\bm{W}_{\bm{z}}\odot\left(\bm{z}^{\prime}_{k}-\frac{1}{\bm{Q}}\odot\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)\right)\,$,
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\bm{I}_{h}^{H}\left(\frac{\bm{W}_{\bm{z}}}{|\bm{Q}|^{2}}\odot\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,$,
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\bm{I}_{h}^{H}\left(\frac{1}{\bm{I}_{H}^{h}\bm{I}_{h}^{H}|\bm{Q}|^{2}}\odot\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,$,
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\bm{I}_{h}^{H}\left(\bm{I}_{H}^{h}\left(\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right)\odot\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,$,
$=-\frac{|\bm{Q}_{H}|^{2}}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\frac{1}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\odot\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$
The proof follows from Proposition H.1 with
$\bm{A}=\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,.$
Appendix I Proof for Proposition 4.4
@@
$\bm{z}_{k}^{+}=\operatornamewithlimits{argmin}_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)+\frac{1}{2}\bm{u}^{\top}\left|\bm{z}_{k}-\bm{z}_{k}^{\prime}\right|^{2}\,$,
@@
$\bm{e}_{k}=\bm{z}_{k}^{+}-\bm{z}_{k}^{\prime}=-\frac{1}{\bm{u}+|\bm{Q}|^{2}}\odot\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\,$,
@@
$\bm{e}_{H,k}=-\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\odot\bm{I}_{h}^{H}\left(\nabla_{\bm{z}_{k}}\widetilde{\Phi}_{k}\left(\bm{z}_{k}^{\prime};\bm{z}_{k}^{(j)}\right)\right)\,.$
@@
$\bm{e}_{k}(\bm{g})=-\frac{1}{\bm{u}+|\bm{Q}|^{2}}\odot\bm{g}\quad\text{and}\quad\bm{e}_{H,k}(\bm{g})=-\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\odot\bm{I}_{h}^{H}\bm{g}\,.$
Since $\|\bm{I}_{h}^{H}\|_{\infty}=\|\bm{I}_{H}^{h}\|_{\infty}=1$ by Lemma E.4, it is straightforward to see that
@@
$\max_{\|\bm{g}\|_{\infty}\leq 1}\left[\left|\bm{e}_{H,k}(\bm{g})\right|\right]_{r}$ $=\max_{\|\bm{g}_{H}\|_{\infty}\leq 1}\left[\left|\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\bm{g}_{H}\right|\right]_{r}\,$,
$=\left[\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\right]_{r}\,$,
@@
$\max_{\|\bm{g}\|_{\infty}\leq 1}\left[\left|\bm{I}_{h}^{H}\bm{e}_{k}(\bm{g})\right|\right]_{r}=\left[\bm{I}_{h}^{H}\frac{1}{\bm{u}+|\bm{Q}|^{2}}\right]_{r}\,.$
Finally, using Lemma E.1, Lemma E.2, and Lemma E.3, we have:
$\bm{u}_{H}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\odot\bm{I}_{h}^{H}\bm{u}\,$,
$\bm{u}_{H}+|\bm{Q}_{H}|^{2}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\odot\bm{I}_{h}^{H}\bm{u}+|\bm{Q}_{H}|^{2}\,$,
$\bm{u}_{H}+|\bm{Q}_{H}|^{2}$ $=\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\odot\bm{I}_{h}^{H}\left(\bm{u}+|\bm{Q}|^{2}\right)\,$,
$\bm{I}_{h}^{H}\left(\frac{1}{\bm{u}+|\bm{Q}|^{2}}\right)$ $=\frac{1}{\bm{u}_{H}+|\bm{Q}_{H}|^{2}}\odot\frac{|\bm{Q}_{H}|^{2}}{\bm{I}_{h}^{H}|\bm{Q}|^{2}}\,$,
$\max_{\|\bm{g}_{2}\|_{\infty}\leq 1}\left[\left|\bm{I}_{h}^{H}\bm{e}_{k}(\bm{g}_{2})\right|\right]_{r}$ $=\max_{\|\bm{g}_{1}\|_{\infty}\leq 1}\left[\left|\bm{e}_{H,k}(\bm{g}_{1})\right|\right]_{r}\quad\text{for all }r\,.$